the value of the expression x3+xy/3+7/2y2 at x=2 y=3 is
 

Grade:6

4 Answers

Priyansh
6 years ago
I believe you wanted to write 3/X not x3 and not 7/2y2 but 7/2y + 2
 
Plugging in the values as x=2 and y=3 
We get
 
2/3 +6/3 +7/6 +2
= 8/3 +7/6 +2
=16/6+7/6+12/6
=16+7+12/6
=35/6
Which is approximately 5.83
Namita kumari
6 years ago
put the value of x and y 
where, x = 2 and y = 3
we get,
x3+xy÷3+7÷2y2
​​8+2×3÷3+7÷12 
8+3+0.58
11.58
therefore the anser of this question is 11.58
 
Preetham Reddy
5 years ago
I think u meant x^3+xy/3 +7/2y^2Putting the values of x and y as 2 and 32^3+2*3/3+7/2*98+2+7/1810+7/18187/7
Sayantan Garai
5 years ago
Answer:
2³+6/3+7/2×3²
=8+2+7/18
=187/18 (Answer).
-----------------------------------------------------